OpenStudy (anonymous):

REWARD FOR HELPING Solve for x: x-6 over 4 =5 14 15 26 8

OpenStudy (jkristia):

\[\frac{x-6}{4} = 5\] start by multiply both sides by 4 to get rid of the fraction.

OpenStudy (anonymous):

Would I multiply everything by 4?

OpenStudy (jkristia):

yes - you always have to do the same on both sides of the equal sign, so \[(4) * (\frac{x-6}{4}) = 5 * 4\] now you can cancel the 4 on the left side, and you are left with x - 6 = ....

OpenStudy (anonymous):

x-6=20 x=24?

OpenStudy (anonymous):

26*

OpenStudy (jkristia):

no x - 6 = 20, what do you do on the left side to 'get rid' of the (-6) ?

OpenStudy (jkristia):

yes

OpenStudy (anonymous):

kay, thanks!